An air bubble has a volume of 1.20 cm^3 when it is
Answer #1

For a definite quantity of ideal gas in transformation, the quantity pV/T is invariant:
pV/T = constant, or
p1V1/T1 = p2V2/T2
V2 = V1(p1/p2)(T2/T1)

V1 = 1.20 cm^3
T2 = T1

p2 = 1 atm = 760 mm Hg
p1 = 760 mm Hg + 35000 mm H2O
the density of mercury with respect to H2O is 13.6
Therefore 35000 mm H2O = 35000/13.6 = 2574 mmHg
p1 = 760 + 2574 = 3334 mm Hg

V2 = V1(p1/p2)(T2/T1) = 1.20×(3334/760)×(1)
V2 = = 1.20 x 3.6 = 4.32 cm^3